I need help with this question.
Given the equation x=-y+9, use substitution to solve 3x-5y=59 for x.
I am working with literal equations.Please help

Answer by stanbon(75887)

Given the equation x=-y+9,
use substitution to solve 3x-5y=59 for x.
---
Substitute for "x" and solve for "y":
3(-y+9)-5y = 59
-3y + 27 -5y = 59
-8y = 22
y = -11/4
-----
Now, solve for "x":
x = -y+9
x = (11/4)+(36/4)
x = 47/4
====================
Cheers,
Stan H.
============